T2 A Ti 2 In time t, an amount of heat O flows through the solid door of area A and thickness d represented above. The temperatures on each side of the door are T2 and Ti, respectively. What factors can be changed that would be certain to decrease O? 3.
Answer #1

Q *t = kA*(T2-T1) / d

where (T2-T1) = temperature difference

d = thickness

A = cross-sectional area

Q can be decreased by decreasing the cross sectional area

increasing the thickness and decreasing the temperature difference